OpenStudy (anonymous):

prove that; sin3(theta)=3sin(theta)-4sin^3(theta)

OpenStudy (anonymous):

sin (1+2)theta

OpenStudy (whpalmer4):

Is that supposed to be \[\sin(3\theta) = 3\sin(\theta)-4\sin^3(\theta)\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sin(1+2)theta=sin1theta *cos2theta+sin2theta*cos1theta

OpenStudy (whpalmer4):

Off the cuff, I would try something with the sum-difference formulas: \[\sin(u\pm v) = \sin u \cos v \pm \cos u \sin v\] You could use \(u = 2\theta\) and \(v = \theta\)

OpenStudy (whpalmer4):

Also \(\cos(2u) = 2\cos^2u-1\)

OpenStudy (whpalmer4):

or \(cos(2u) = 1-2\sin^2u\)

OpenStudy (anonymous):

=sin1theta(cos^2theta-sin^2theta)+(2sin1theta*cos1 theta)*cos 1 theta

OpenStudy (anonymous):

ok what i dont understand is how sin 3(theta) can equal sin2(theta)cos(theta) +cos2(theta)sin(theta) and then = 2sin(theta)cos(theta)cos(theta)+(1-2(sin^2theta)sin(theta)?

OpenStudy (anonymous):

u just sin 3 theta=sin (2+1)theta

OpenStudy (anonymous):

sin(1+2)theta=sin1theta *cos2theta+sin2theta*cos1theta

OpenStudy (anonymous):

ok i got that but the (1-2sin^2theta) bit is what i dont understand in that part

OpenStudy (anonymous):

use cos2theta=cos^2theta-sin^2theta

OpenStudy (anonymous):

sin^2theta=2sin1theta*costheta